What is an equation of the line that is perpendicular to 2y+2= -7/4(x-7) and contains (-2,-3)?
thanks if you can help

Answer by stanbon(75887)

What is an equation of the line that is perpendicular to 2y+2= -7/4(x-7) and contains (-2,-3)?
-----
Find the slope of the given equation:
2y = (-7/4)x + (49/4) - 2
y = (-7/8)x + some constant
---
Therefore its slope is -7/8
-----
Any line perpendicular to it must have slope = 8/7
----
Your Equation:
Form: y = mx + b
You know m = 8/7 and y = -3 when x = -2
Solve for "b":
-3 = (8/7)(-2)+ b
-3 = -16/7 + b
b = (-21/7)+(16/7)
b = -5/7
-----
Equation:
y = (8/7)x - (5/7)
====================
Cheers,
Stan H.
=============